OpenStudy (anonymous):

m-4n=8, solve for m please explain

OpenStudy (anonymous):

m=4n+8

OpenStudy (anonymous):

i came up with m=8-4n what did i do wrong?

jimthompson5910 (jim_thompson5910):

you need to add 4n to both sides

jimthompson5910 (jim_thompson5910):

adding 4n to -4n on the left side gives you -4n+4n = 0n = 0 which goes away

OpenStudy (anonymous):

oh! i subtracted 4n from each side

jimthompson5910 (jim_thompson5910):

adding 4n to the right side gives you 8+4n

jimthompson5910 (jim_thompson5910):

so your answer is m = 8+4n which can be written as m = 4n+8

OpenStudy (anonymous):

ok. i think i did -4n instead of +4n

jimthompson5910 (jim_thompson5910):

yep or you just moved the -4n term away from what you wanted to solve for

OpenStudy (anonymous):

can i ask you another one then? st+3t=6, solve for s

jimthompson5910 (jim_thompson5910):

ok follow the order of operations in reverse So you're undoing addition/subtraction

myininaya (myininaya):

what happens if you did subtract 4n instead of adding 4n you would get the following: m-4n-4n=8-4n m-8n=8-4n <-- this made things uglier

jimthompson5910 (jim_thompson5910):

so you're undoing +3t, so you subtract 3t from both sides to get st = 6-3t

OpenStudy (anonymous):

ok i understand i did it wrong thank you very much

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what comes next?

OpenStudy (anonymous):

s=6-3t/t, is that right?

OpenStudy (anonymous):